OpenStudy (mathwizz):

i need to solve 30= 6 + 4x

OpenStudy (amistre64):

yes you do :) what should we do first?

OpenStudy (mathwizz):

isolate the x

OpenStudy (amistre64):

good; so first we get rid of that 6; subtract 6 from both sides of the = sign like this: 30-6 = 6-6+4x 24 = 4x ; now what? how do we undo a multiplication of 4x?

OpenStudy (anonymous):

yep

OpenStudy (mathwizz):

do u divide 4 by 24

OpenStudy (amistre64):

thats a good thought, but a little backwards; lets divide everything by 4 tho, since that is stuck to the x 24/4 = 4x/4 6 = 1x 6 = x

OpenStudy (amistre64):

very good :)

OpenStudy (mathwizz):

thanks

OpenStudy (mathwizz):

ive got another one

OpenStudy (amistre64):

if you post it up again, someone else may be able to help you, th elibrary is closing soon and I gotta go..

OpenStudy (anonymous):

dang it amistre... u keep getting to all of them first! >.<